OpenStudy (anonymous):

I dont understand how you get from one step to another. . . has to do with differential equasions

OpenStudy (anonymous):

\[u-70=e ^{\ln(u-70)} = e ^{kt+C} = e ^{C}e ^{kt}\]

OpenStudy (anonymous):

ln(u-70) + kt+C

OpenStudy (anonymous):

theyre in reverse order

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

u-7 = u-70 well it simpliifies to this. by using the exponent properties right?

OpenStudy (anonymous):

i think. . . i'm not sure

OpenStudy (anonymous):

I am sure :)